OpenStudy (anonymous):

Change the radical to simplest radical form. 5square 50

OpenStudy (anonymous):

Change the radical to simplest radical form.\[5\sqrt{50}\]

jhonyy9 (jhonyy9):

rewrite this 5sqrt(2*25) =

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[25\sqrt{5}\] The answer right?

jhonyy9 (jhonyy9):

not is right !!!

jhonyy9 (jhonyy9):

sqrt50 =sqrt(2*25) =sqrt(2*5^2) =5sqrt2

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so than good luck bye was my pleasure

OpenStudy (anonymous):

Ok, Thanks